OpenStudy (anonymous):

Find the speed parameter of a particle that takes 2.8 y longer than light to travel a distance of 7.7 ly.

OpenStudy (anonymous):

Please define the speed parameter. Do you mean the Lorentz factor? If so, the Lorentz factor is defined with respect to temporal dilation as:\[\frac{ \Delta t' }{ \Delta t }=\gamma =\frac{ 1 }{ \sqrt{1-\frac{ v ^{2} }{ c ^{2} }} }\]where ∆t' is the time in the moving frame; ∆t is the time in the rest frame; v is the velocity; and c is the speed of light.